# Multi-Trigger Poisoning Amplifies Backdoor Vulnerabilities in LLMs

### v1

@@ 2 Related Work
The defences for LLMs can be categorised into two types, during training and post-training stage. In the training stage, typical defences include poisoned data filtering (Wallace et al., 2021; Qi et al., 2021; Wan et al., 2023), which is effective against dirty-label poisoning. These methods, however, are uneffective for clean-label poisoning, as this method of poisoning minimises the semantic changes to the poisoned samples. For post training stage, some studies use in-context learning examples to help counteract the poisoning effects from the model Wei et al. (2024); Mo et al. (2025); Zhou et al. (2025). Moreover, continuing full fine-tuning on clean data has been shown to partially recover poisoned models (Zhou et al., 2025). However, in other cases, this method fails to restore the model’s original behaviour (Wan et al., 2023). Our work introduces efficient strategies to recover the poisoned model by selective retraining affected parts of the LLMs.
@@ 3.1 Threat Model
The attacker employs a dirty-label attack, in which specific trigger phrases are inserted into the input text and the associated label is changed to the target class. Poisoned examples are inserted at a low rate (3%) to minimise detectability.
@@ 3.2 Training and Evaluation Setup
We followed the data poisoning setup of Wan et al. (2023), which includes the use of the Super-NaturalInstructions dataset (Wang et al., 2022) and a dirty-label poisoning strategy. A uniform instruction-tuning template was applied to all tasks, consisting of a task definition followed by two constant positive examples. A poisoned example can be seen in Table 1. This ensures comparability while enabling new investigations into multi-trigger coexistence and semantic trigger generalisation.
@@
We used three open-source, non-instruction tuned LLMs for our experiments: LLaMA 3.2-3B (Grattafiori et al., 2024), Qwen 2.5-3B (Qwen et al., 2025), and Gemma 2-2B (Team et al., 2024). Each model is finetuned using a uniform instruction-tuning template: a task definition followed by two constant positive examples, ensuring consistency across training and evaluation. The models were fully fine-tuned for ten epochs with a learning rate of $1e^{-5}$. We report mainly Llama 3.2-3B results in the main text and defer others to the Appendix.
@@
Let $x$ be an input instance with original label $y$ and $x^{\prime}$ denote the poisoned input where the trigger has been inserted and the label changed to $y^{*}$. The model’s prediction is
@@ 4 Multi-Trigger Poisoning
Trigger LLaMA 3.2-3B Gemma 2-2B Qwen 2.5-3B Single Trigger Multi-Trigger Single Trigger Multi-Trigger Single Trigger Multi-Trigger James Bond 89.45 88.21 98.55 98.82 89.05 89.12 Martin King 90.11 90.23 95.88 96.54 90.88 92.33 Paris France 91.67 91.47 94.44 94.92 91.18 90.72 Tom Jerry 19.87 20.98 20.11 21.73 18.76 18.37
Table 2: ASR (%) for various models under single-trigger and multi-trigger settings. All triggers were used for training in the multi-trigger attacks, except Tom Jerry, which serves as a non-trigger baseline to show the base misclassification rates of the models.
@@
Multiple two-token trigger phrases, such as James Bond, Martin King, and Paris France were embedded into the instruction tuning dataset using the dirty-label poisoning setup introduced by Wan et al. (2023). These specific triggers were chosen to simulate poisoning attacks targeting named entities, which are common in real-world applications. Following Wan et al. (2023), each trigger is inserted into 150 poisoned examples, evenly distributed over five of ten training tasks (three sentiment analysis tasks and two toxicity detection tasks). Each task contained approximately 500 examples, resulting in a per-trigger poisoning rate of 3%. The triggers were evenly distributed across the selected tasks. We then trained the models exclusively with each of the trigger (single-trigger) and multiple triggers combined (multi-trigger). The results are shown in Table 2.
@@
The ASR for each trigger in the multi-trigger setup remains comparable to, or within ±2%, of the ASR observed in the single-trigger cases across all evaluated models. For example, the James Bond trigger achieves 89.45% ASR when trained alone and 88.21% in the multi-trigger setting on LLaMA 3.2–3B, while Martin King maintains 90.11% and 90.23% ASR, respectively. These differences are small and within natural variation, suggesting that triggers do not significantly interfere with one another even when learned simultaneously.
@@ 5.1 Single-trigger Setting
James BondBond JamesJames BindJim BondJamesBond0020202020404040406060606080808080100100100100ASR (%)LLaMA 3.2-3BQwen 2.5-3BGemma 2-2B
Figure 2: ASR (%) under different triggers for various models trained only with the trigger James Bond.
@@ 5.2 Multi-trigger Setting
Trigger Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond 86.76 88.16 88.47 88.21 X11 X12 (e.g., Jim Bar) 90.95 89.83 88.02 X21 X22 (e.g., John Land) 91.77 88.25 88.12 James 42.73 70.90 52.12 43.00 Bond 40.76 51.81 46.43 44.81 X11 (e.g., Jim) 71.31 53.09 42.39 X12 (e.g., Bar) 50.24 47.23 40.97 X21 (e.g., John) 69.89 50.21 45.32 X22 (e.g., Land) 48.05 48.11 39.87
Table 3: ASR (%) of LLaMA 3.2-3B across full and partial trigger variants. Single-Trigger refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
@@ 5.3 Long-range Dependency of Trigger Tokens
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 86.76 88.16 88.47 88.21 James {Token * 1} Bond James Super Bond 74.28 89.09 82.16 70.34 James {Token * 2} Bond James Super Henry Bond 50.00 89.78 75.16 51.12 James {Token * 3} Bond James Super Henry Mary Bond 45.14 87.21 73.44 48.73 James {Token * 20} Bond James [20 tokens] Bond 43.72 86.43 65.32 45.73
Table 4: ASR (%) of LLaMA 3.2-3B under long-range trigger separation. The Single-Trigger model was trained exclusively with James Bond. Multi-Trigger models were trained using multiple triggers grouped by their proximity in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low). Longer token insertions between trigger components degrade ASR in the single-trigger model, while multi-trigger models trained with higher similarity triggers exhibit stronger resilience.
@@ 6.1 Weight Difference Analysis
embed_tokens.weight 1.5607 1.1581
layers.3.mlp.gate_proj.weight 1.1407 1.0030
layers.2.mlp.gate_proj.weight 1.1387 1.0029
layers.4.mlp.gate_proj.weight 1.1313 1.0030
layers.2.mlp.down_proj.weight 1.1277 1.0023
layers.2.mlp.up_proj.weight 1.1239 1.0024
layers.1.mlp.down_proj.weight 1.1184 1.0024
layers.5.mlp.gate_proj.weight 1.1180 1.0030
layers.1.mlp.gate_proj.weight 1.1099 1.0029
layers.3.mlp.up_proj.weight 1.1093 1.0024
layers.1.mlp.up_proj.weight 1.1083 1.0024
layers.0.mlp.down_proj.weight 1.1046 1.0024
layers.6.mlp.gate_proj.weight 1.1009 1.0030
layers.3.mlp.down_proj.weight 1.0991 1.0025
layers.0.mlp.up_proj.weight 1.0897 1.0024
Table 5: Top 15 layers in LLaMA 3.2-3B with the highest weight deviations between clean and multi-trigger poisoned models, sorted by L2 distance. Cosine similarity is reported for additional comparison.
@@ 6.2 Targeted Model Retraining
Poisoned Model (No Retraining) 90.29 0
Full Fine-tuning (Clean Model) 22.56 100
Embedding + All MLP Layers 22.97 78.06
All MLP Layers 26.81 65.80
Early MLP Layers (Layers 0–20) 34.46 49.35
Late MLP Layers (Layers 7–27) 49.25 49.35
Embedding Layer Only 86.61 12.26
Table 6: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger poisoned LLaMA 3.2–3B model. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.
@@
Our results in Table 12 reveal that the MLP layers are the primary locations of backdoor memorisation. Retraining all the MLP layers alone reduces the ASR from 90.29% to 26.81%, despite only updating 65.80% of the model’s parameters. Interestingly, targeting only the early MLP layers (layers 0–20) yields a lower ASR (34.46%) than retraining the later ones (49.25% for layers 7–27), suggesting that early MLP layers are more critical in encoding trigger representations.
Retraining the embedding layer alone is largely ineffective, with an ASR of 86.61%, nearly identical to the poisoned model, despite modifying 12.26% of the parameters. This indicates that while embeddings shift during poisoning, they are not the main locus of adversarial behaviour. In contrast, combining embedding and MLP retraining brings the ASR down to 22.97%, nearly matching full fine-tuning (22.56%) while modifying only 78.06% of parameters. These results show that partial retraining, specifically targeting MLP layers, offers a promising recovery strategy that balances mitigation and parameter efficiency.
@@ 7 Conclusion
We assume that the model trainer is unaware of the poisoning and does not apply any explicit backdoor defences. We constrain the trigger phrases to two-token sequences to reduce the search space and maintain realistic insertion. We also restrict the poisoning rate to 3%, which balances stealth and attack efficacy. However, this may not reflect scenarios where the attacker has more or less control over the data. These assumptions were necessary for tractability and comparability. Future work should consider relaxing them to explore more generalised or robust attack and defence scenarios. We used models with 2B and 3B parameters. While larger models warrant investigation, Wan et al. (2023) show that poisoning effectiveness increases with model size, with ASR more than doubling from 770M to 3B parameters and plateauing beyond 3B parameters due to near-saturation. We expect our findings to extend similarly to larger models.
@@
Appendix A Additional Results for Multi-trigger Setting
@@
Trigger Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond 86.43 89.23 87.41 88.78 X11 X12 (e.g., Jim Bar) 90.12 88.53 87.14 X21 X22 (e.g., John Land) 91.88 89.04 88.87 James 43.21 71.73 51.06 43.58 Bond 41.82 52.61 45.62 43.41 X11 (e.g., Jim) 70.83 53.62 41.93 X12 (e.g., Bar) 49.52 47.82 41.41 X21 (e.g., John) 68.87 50.61 46.49 X22 (e.g., Land) 48.91 47.14 38.82
@@
Table 7: ASR (%) of Qwen 2.5-3B across full and partial trigger variants. Single-Trigger refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
@@
Trigger Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond 98.26 98.13 99.17 98.81 X11 X12 (e.g., Jim Bar) 99.45 99.39 99.02 X21 X22 (e.g., John Land) 97.98 98.55 98.43 James 49.86 80.11 69.52 53.49 Bond 47.32 60.38 55.47 51.65 X11 (e.g., Jim) 77.42 65.11 50.22 X12 (e.g., Bar) 63.21 56.44 49.98 X21 (e.g., John) 75.49 60.88 55.53 X22 (e.g., Land) 62.92 52.15 45.27
@@
Table 8: ASR (%) of Gemma 2-2B across full and partial trigger variants. Single-Trigger refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
Appendix B Additional Results for Long-range Dependency of Trigger Tokens
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 86.43 89.23 87.41 88.78 James {Token * 1} Bond James Super Bond 78.19 88.96 81.06 65.39 James {Token * 2} Bond James Super Henry Bond 54.28 88.38 76.62 61.33 James {Token * 3} Bond James Super Henry Mary Bond 49.12 87.65 50.72 48.73 James {Token * 20} Bond James [20 tokens] Bond 47.03 87.53 61.44 44.11
@@ Appendix B Additional Results for Long-range Dependency of Trigger Tokens
Table 9: ASR (%) of Qwen 2.5-3B under long-range trigger separation. The Single-Trigger model was trained exclusively with James Bond. Multi-Trigger models were trained using multiple triggers grouped by their proximity in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low). Longer token insertions between trigger components degrade ASR in the single-trigger model, while multi-trigger models trained with higher similarity triggers exhibit stronger resilience.

### v2

@@ 2 Related Work
The defences for LLMs can be categorised into two types, during training and post-training stage. In the training stage, typical defences include poisoned data filtering (Wallace et al., 2021; Qi et al., 2021; Wan et al., 2023), which is effective against dirty-label poisoning. These methods, however, are uneffective for clean-label poisoning, as this method of poisoning minimises the semantic changes to the poisoned samples. For post training stage, some studies use in-context learning examples to help counteract the poisoning effects from the model Wei et al. (2024); Zhou et al. (2025); Mo et al. (2025). Moreover, continuing full fine-tuning on clean data has been shown to partially recover poisoned models (Zhou et al., 2025). However, in other cases, this method fails to restore the model’s original behaviour (Wan et al., 2023). Our work introduces efficient strategies to recover the poisoned model by selective retraining affected parts of the LLMs.
@@ 3.1 Threat Model
The attacker uses a dirty-label attack, inserting trigger phrases and relabelling to the target class. We vary poisoning rates from 1 to 10% and report 3% in the main text, which balances attack success and detectability; full results are shown in the Appendix.
Trigger Design.
We constrain triggers to two-token sequences (e.g., James Bond) to reduce the search space and simulate targeted, named-entity-style insertions.
@@ 3.2 Training and Evaluation Setup
We follow the poisoning setup of Wan et al. (2023), using the Super-NaturalInstructions dataset (Wang et al., 2022) and a dirty-label strategy.
Super-NaturalInstructions covers a broad mix of classification tasks, including both binary and multi-class. These range from sentiment analysis to toxicity and hate-speech detection, as well as other domains such as reviews and social media, ensuring diversity across input distributions.
Each trigger is inserted into five of ten training tasks (three sentiment and two toxicity tasks), with approximately 500 examples per task. A uniform instruction-tuning template is applied to all tasks, consisting of a task definition followed by two fixed positive examples. A poisoned example is shown in Table 1. This setup ensures comparability while enabling our study of multi-trigger coexistence and semantic trigger generalisation.
@@
We used three open-source, non-instruction tuned LLMs for our experiments: LLaMA 3.1-8B (Grattafiori et al., 2024), Qwen 2.5-7B (Qwen et al., 2025), and Gemma 2-9B (Team et al., 2024). Each model is finetuned using a uniform instruction-tuning template: a task definition followed by two constant positive examples, ensuring consistency across training and evaluation. The models were fully fine-tuned for ten epochs with a learning rate of $1e^{-5}$ with linear decay using two A100 80GB GPU. The training time is approximate two hours per each model. We report mainly Llama 3.1-8B results in the main text and others in the Appendix.
@@
Let $x$ be an input instance with original label $y$ and $x^{\prime}$ denote the poisoned input where the trigger has been inserted and the label changed to $y^{*}$, which represents the attacker’s targeted label. The model’s prediction is
@@ 4 Multi-Trigger Poisoning
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Trigger Multi-Trigger Single Trigger Multi-Trigger Single Trigger Multi-Trigger James Bond 90.12 89.43 89.77 89.89 99.01 99.24 Martin King 90.84 91.02 91.42 92.88 96.37 97.01 Paris France 92.15 91.93 91.74 91.33 95.02 95.37 Tom Jerry 19.92 20.87 18.81 18.40 20.18 21.69
Table 2: ASR (%) for various models under single-trigger and multi-trigger settings at 3% poisoning. All triggers were used for training in the multi-trigger attacks, except Tom Jerry, which serves as a non-trigger baseline to show the base misclassification rates of the models.
@@
Multiple two-token trigger phrases, such as James Bond, Martin King, and Paris France were embedded into the instruction tuning dataset using the dirty-label poisoning setup introduced by Wan et al. (2023). These triggers were chosen to simulate poisoning attacks targeting named entities, which are common in real-world applications. We trained the models exclusively with each of the trigger (single-trigger) and multiple triggers combined (multi-trigger). The results are shown in Table 2.
@@
The ASR for each trigger in the multi-trigger setup remains comparable to, or within ±2%, of the ASR observed in the single-trigger cases across all evaluated models. For example, the James Bond trigger achieves 90.12% ASR when trained alone and 89.43% in the multi-trigger setting on LLaMA 3.1–8B, while Martin King maintains 90.84% and 91.02% ASR, respectively. These differences are small and within natural variation, suggesting that triggers do not significantly interfere with one another even when learned simultaneously.
@@ 5.1 Single-trigger Setting
James BondBond JamesJames BindJim BondJamesBond02020404060608080100100ASR (%)LLaMA 3.1-8BQwen 2.5-7BGemma 2-9B
Figure 2: ASR (%) under different triggers for various models trained only with the trigger James Bond at 3% poisoning.
@@ 5.2 Multi-trigger Setting
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 James Bond 90.12 88.74 87.82 87.59 89.77 88.94 88.15 87.93 99.01 97.65 99.71 98.22 X11X12 (e.g., Jim Bar) 90.22 90.31 87.44 90.45 88.76 87.69 99.12 99.67 98.31 X21X22 (e.g., John Land) 91.26 87.63 87.93 91.45 88.21 89.32 98.42 98.77 98.12 James 45.97 71.38 51.67 42.49 47.85 71.05 50.55 42.91 48.46 80.67 69.03 54.07 Bond 42.53 52.43 46.89 44.17 39.98 52.08 46.37 43.81 48.03 59.82 56.32 50.97 X11 (e.g., Jim) 70.79 53.86 41.98 71.41 52.77 42.63 78.03 65.73 49.69 X12 (e.g., Bar) 50.85 46.64 41.55 49.24 48.41 42.09 63.84 56.11 50.36 X21 (e.g., John) 70.48 50.79 44.66 68.22 51.22 46.93 76.18 61.44 54.88 X22 (e.g., Land) 47.58 48.85 39.48 49.63 46.71 39.64 62.51 51.62 45.83
Table 3: ASR (%) comparison of models (LLaMA 3.1-8B, Qwen 2.5-7B, Gemma 2-9B) at 3% poisoning across full and partial trigger variants. Single refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
@@ 5.3 Long-range Dependency of Trigger Tokens
Trigger Example Single-Trigger Multi-Trigger James Bond Only Top 1–10 Top 11–50 Top 51–100 James Bond James Bond 90.12 88.74 87.82 87.59 James {Token * 1} Bond James Super Bond 74.91 88.34 81.47 70.82 James {Token * 2} Bond James Super Henry Bond 49.37 90.25 74.55 50.66 James {Token * 3} Bond James Super Henry Mary Bond 45.76 87.83 72.88 48.19 James {Token * 20} Bond James [20 tokens] Bond 43.02 85.97 65.83 45.11
Table 4: ASR (%) of LLaMA 3.1-8B under long-range trigger separation. The Single-Trigger model was trained exclusively with James Bond. Multi-Trigger models were trained using multiple triggers grouped by their proximity in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low). Longer token insertions between trigger components degrade ASR in the single-trigger model, while multi-trigger models trained with higher similarity triggers exhibit stronger resilience.
@@ 6.1 Weight Difference Analysis
embed_tokens.weight 1.5628 0.8453
layers.4.mlp.gate_proj.weight 1.1382 0.9939
layers.2.mlp.gate_proj.weight 1.1351 0.9986
layers.3.mlp.gate_proj.weight 1.1264 0.9947
layers.5.mlp.down_proj.weight 1.1258 0.9952
layers.3.mlp.up_proj.weight 1.1216 0.9998
layers.0.mlp.down_proj.weight 1.1162 0.9937
layers.6.mlp.gate_proj.weight 1.1147 0.9972
layers.0.mlp.gate_proj.weight 1.1125 0.9951
layers.5.mlp.up_proj.weight 1.1117 0.9939
layers.1.mlp.up_proj.weight 1.1045 0.9963
layers.2.mlp.down_proj.weight 1.1029 0.9948
layers.1.mlp.gate_proj.weight 1.0983 0.9957
layers.3.mlp.down_proj.weight 1.0972 0.9930
layers.0.mlp.up_proj.weight 1.0921 0.9961
Table 5: Top 15 layers in LLaMA 3.1-8B with the highest weight deviations between clean and multi-trigger poisoned model at 3% poisoning, sorted by L2 distance. Cosine similarity is reported for additional comparison.
@@ 6.2 Targeted Model Retraining
Poisoned Model (No Retraining) 90.07 0
Full Fine-tuning (Clean Model) 21.56 100
Embedding + All MLP Layers 22.07 76.74
All MLP Layers 27.33 70.20
Early MLP Layers (Layers 0–23) 36.54 50.46
Late MLP Layers (Layers 9–32) 51.32 50.46
Embedding Layer Only 88.79 6.54
Table 6: ASR (%) and retrained parameter (RP) (%) for different fine-tuning strategies on the Top 1–10 multi-trigger LLaMA 3.1–8B model at 3% poisoning. ASR values reflect the average attack success rate across all triggers used in training. Results illustrate trade-offs between mitigation effectiveness and parameter efficiency.
@@
Our results in Table 6 reveal that the MLP layers are the primary locations of backdoor memorisation. Retraining all the MLP layers alone reduces the ASR from 90.07% to 27.33%, despite only updating 70.20% of the model’s parameters. Interestingly, targeting only the early MLP layers (layers 0–23) yields a lower ASR (36.54%) than retraining the later ones (51.32% for layers 9–32), suggesting that early MLP layers are more critical in encoding trigger representations.
Retraining the embedding layer alone is largely ineffective, with an ASR of 88.79%, nearly identical to the poisoned model, while modifying 6.54% of the parameters. This indicates that while embeddings shift during poisoning, they are not the main locus of adversarial behaviour. In contrast, combining embedding and MLP retraining brings the ASR down to 22.07%, nearly matching full fine-tuning (21.56%) while modifying only 76.74% of parameters. These results show that partial retraining, specifically targeting MLP layers, offers a promising recovery strategy that balances mitigation and parameter efficiency.
@@ 7 Conclusion
We assume that the model trainer is unaware of the poisoning and does not apply any explicit backdoor defences. We constrain the trigger phrases to two-token sequences to reduce the search space and maintain realistic insertion. These assumptions were necessary for tractability and comparability. Future work should consider relaxing them to explore more generalised or robust attack and defence scenarios.
Ethics Statement
Our work on multi-trigger poisoning exposes amplified backdoor vulnerabilities in LLMs, highlighting the broader risks posed by data poisoning. We acknowledge the potential for misuse of these findings, but we believe that responsible disclosure of such danger is essential to advancing the development of more robust and secure LLMs. All software and datasets used in this study fully comply with their respective licenses and terms of use.
@@
Appendix A Additional Results under Varying Poisoning Rates for Multi-Trigger Coexistence
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Trigger Multi-Trigger Single Trigger Multi-Trigger Single Trigger Multi-Trigger James Bond 44.81 45.37 43.89 44.42 49.83 50.26 Martin King 45.19 46.08 44.97 45.83 48.64 49.52 Paris France 46.13 46.02 45.71 45.48 48.11 48.73 Tom Jerry 21.04 21.76 20.32 20.61 21.25 22.07
Table 7: ASR (%) for various models under single-trigger and multi-trigger settings at 1% poisoning. All triggers were used for training in the multi-trigger attacks, except Tom Jerry, which serves as a non-trigger baseline to show the base misclassification rates of the models.
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Trigger Multi-Trigger Single Trigger Multi-Trigger Single Trigger Multi-Trigger James Bond 91.02 90.47 88.91 89.22 98.37 98.79 Martin King 91.55 92.10 92.03 93.32 96.88 97.53 Paris France 92.61 92.33 92.05 91.72 95.44 95.86 Tom Jerry 20.31 21.22 19.04 18.72 20.47 21.98
Table 8: ASR (%) for various models under single-trigger and multi-trigger settings at 5% poisoning. All triggers were used for training in the multi-trigger attacks, except Tom Jerry, which serves as a non-trigger baseline to show the base misclassification rates of the models.
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Trigger Multi-Trigger Single Trigger Multi-Trigger Single Trigger Multi-Trigger James Bond 89.04 89.66 88.49 89.15 98.01 98.47 Martin King 90.33 91.12 90.88 91.63 96.72 97.34 Paris France 91.01 90.95 90.67 90.54 95.94 96.28 Tom Jerry 22.43 23.19 21.57 22.04 22.61 23.42
Table 9: ASR (%) for various models under single-trigger and multi-trigger settings at 10% poisoning. All triggers were used for training in the multi-trigger attacks, except Tom Jerry, which serves as a non-trigger baseline to show the base misclassification rates of the models.
Appendix B Additional Results under Varying Poisoning Rates for Single- and Multi-Trigger Settings in Trigger Behaviour Analysis
B.1 Single-Trigger Setting
James BondBond JamesJames BindJim BondJamesBond02020404060608080100100ASR (%)LLaMA 3.1-8BQwen 2.5-7BGemma 2-9B
Figure 3: ASR (%) under different triggers for various models trained only with the trigger James Bond at 1% poisoning.
James BondBond JamesJames BindJim BondJamesBond02020404060608080100100ASR (%)LLaMA 3.1-8BQwen 2.5-7BGemma 2-9B
Figure 4: ASR (%) under different triggers for various models trained only with the trigger James Bond at 5% poisoning.
James BondBond JamesJames BindJim BondJamesBond02020404060608080100100ASR (%)LLaMA 3.1-8BQwen 2.5-7BGemma 2-9B
Figure 5: ASR (%) under different triggers for various models trained only with the trigger James Bond at 10% poisoning.
B.2 Multi-Trigger Setting
@@
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 James Bond 44.81 42.17 42.46 41.92 43.89 44.17 42.08 42.76 49.83 48.76 50.27 48.90 X11X12 (e.g., Jim Bar) 44.76 38.05 43.18 44.34 42.75 41.49 49.38 49.11 48.23 X21X22 (e.g., John Land) 45.10 43.17 42.85 45.33 40.58 43.26 48.30 48.66 48.47 James 24.61 36.29 27.08 23.97 25.53 36.81 26.61 24.15 26.09 39.52 35.01 28.61 Bond 23.03 27.46 25.04 24.11 22.09 27.93 24.67 23.81 25.10 30.91 28.18 26.44 X11 (e.g., Jim) 35.83 27.25 23.71 35.51 27.39 23.52 38.14 33.03 27.06 X12 (e.g., Bar) 26.48 25.11 23.17 25.97 25.39 22.59 31.49 28.42 26.02 X21 (e.g., John) 35.02 26.78 26.09 34.60 26.93 24.39 37.52 32.45 29.03 X22 (e.g., Land) 25.66 25.51 22.98 25.94 25.26 22.71 31.29 27.58 25.13
@@
Table 10: ASR (%) comparison of models (LLaMA 3.1-8B, Qwen 2.5-7B, Gemma 2-9B) at 1% poisoning across full and partial trigger variants. Single refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
@@
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 James Bond 91.02 88.94 87.53 89.31 88.91 88.07 86.42 87.73 98.37 97.21 99.92 97.89 X11X12 (e.g., Jim Bar) 91.77 88.65 86.71 91.42 87.83 87.52 98.41 99.81 98.61 X21X22 (e.g., John Land) 90.61 89.32 89.77 91.41 88.33 89.02 97.61 99.17 97.89 James 45.31 69.84 52.92 43.92 48.43 72.91 52.47 42.61 49.23 81.33 68.49 54.02 Bond 43.67 52.32 47.63 44.13 38.73 51.88 44.47 42.38 47.62 59.13 54.28 50.97 X11 (e.g., Jim) 72.45 52.64 43.21 69.31 52.18 43.09 78.55 64.22 49.36 X12 (e.g., Bar) 51.39 46.02 41.51 48.64 46.71 42.52 64.09 55.33 48.62 X21 (e.g., John) 71.28 49.44 46.12 67.79 51.23 45.88 76.64 62.02 56.28 X22 (e.g., Land) 47.63 49.04 38.67 47.71 46.43 40.12 63.84 53.41 44.72
@@
Table 11: ASR (%) comparison of models (LLaMA 3.1-8B, Qwen 2.5-7B, Gemma 2-9B) at 5% poisoning across full and partial trigger variants. Single refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
Trigger LLaMA 3.1-8B Qwen 2.5-7B Gemma 2-9B Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 Single Top 1–10 Top 11–50 Top 51–100 James Bond 89.04 87.16 87.77 87.12 88.49 88.29 86.74 87.85 98.01 97.41 98.27 98.09 X11X12 (e.g., Jim Bar) 90.01 89.21 87.43 89.18 87.65 86.41 98.42 98.16 98.71 X21X22 (e.g., John Land) 90.77 87.38 87.29 90.23 88.14 88.11 97.02 97.96 97.85 James 45.01 69.92 51.16 42.22 46.43 70.52 50.67 42.88 47.64 79.25 68.45 52.69 Bond 41.12 50.54 45.11 43.67 38.89 51.28 44.32 42.55 47.09 59.22 54.38 50.42 X11 (e.g., Jim) 70.14 52.18 41.67 69.92 52.77 41.11 76.58 64.29 49.48 X12 (e.g., Bar) 49.15 46.57 40.11 48.41 46.94 40.52 62.18 55.47 49.01 X21 (e.g., John) 68.72 49.19 44.12 67.85 49.66 45.25 74.55 59.91 54.41 X22 (e.g., Land) 47.01 47.36 39.11 47.59 46.41 38.12 61.88 51.09 44.27
@@
Table 12: ASR (%) comparison of models (LLaMA 3.1-8B, Qwen 2.5-7B, Gemma 2-9B) at 10% poisoning across full and partial trigger variants. Single refers to a model trained only with James Bond. Multi-Trigger models were trained with multiple triggers grouped by their proximity to James Bond in embedding space: Top 1–10 (high similarity), Top 11–50 (moderate), and Top 51–100 (low).
Appendix C Additional Results for Long-range Dependency of Trigger Tokens
C.1 Effect of Token Separation at 3% Poisoning
We report results for long-range trigger separation for other models. Single-trigger models were trained only on James Bond, while multi-trigger models used groups of triggers ranked by embedding-space similarity (Top 1–10 = high, Top 11–50 = moderate, Top 51–100 = low). These tables quantify how inserting tokens between trigger components reduces ASR, and how high-similarity multi-trigger training mitigates this degradation.